OpenStudy (anonymous):

Solve the quadratic equation v2 - 2v = 3 by completing the square. A. {2, -3} B. {3, -1} C. {1, 3} D. {-3, 2}

OpenStudy (anonymous):

@e.cociuba

OpenStudy (nurali):

v² - 2v = 3 v² - v = 3 + (- 1)² v² - v = 3 + 1 (v - 1)² = 4 v - 1 = ± 2 v = 1 + 2, v = 3 v = 1 - 2, v = - 1 Answer: v = 3, - 1; factors: (v - 3)(v + 1)

OpenStudy (e.cociuba):

B.)

OpenStudy (e.cociuba):

It is correct! :)